OpenStudy (anonymous):

Medal Which ratio forms a proportion with 6 over 15? answers below

OpenStudy (anonymous):

\[\frac{ 8 }{ 10 }\] \[\frac{ 5 }{ 2 }\] \[\frac{ 3 }{ 12 }\] \[\frac{ 8 }{ 20 }\]

OpenStudy (anonymous):

@jigglypuff314 can you help me?

jigglypuff314 (jigglypuff314):

simplify 6/15 by dividing top and bottom by 3 so your simplified fraction would be 2/5 2/5 also equals what else?

OpenStudy (anonymous):

8/20

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

4 x 2 = 8 4 x 5 = 20

OpenStudy (anonymous):

thank you!!

jigglypuff314 (jigglypuff314):

your welcome :D